OpenStudy (anonymous):

graph the equation 4x+y=9

OpenStudy (anonymous):

y-intercept form y=mx+b so solce for y y=-4x+9 b=(0,9) starting point and m(slope)=-4/1 so go down 4 and right 1 from (0,9) to get second point connect two points with a line

OpenStudy (anonymous):

and graph the equation 2x-3y=-6